# Test-Time Model Adaptation with Only Forward Passes

### v1

@@ Abstract
Test-time adaptation has proven effective in adapting a given trained model to unseen test samples with potential distribution shifts. However, in real-world scenarios, models are usually deployed on resource-limited devices, e.g., FPGAs, and are often quantized and hard-coded with non-modifiable parameters for acceleration. In light of this, existing methods are often infeasible since they heavily depend on computation-intensive backpropagation for model updating that may be not supported. To address this, we propose a test-time Forward-Only Adaptation (FOA) method. In FOA, we seek to solely learn a newly added prompt (as model’s input) via a derivative-free covariance matrix adaptation evolution strategy. To make this strategy work stably under our online unsupervised setting, we devise a novel fitness function by measuring test-training statistic discrepancy and model prediction entropy. Moreover, we design an activation shifting scheme that directly tunes the model activations for shifted test samples, making them align with the source training domain, thereby further enhancing adaptation performance. Without using any backpropagation and altering model weights, FOA runs on quantized 8-bit ViT outperforms gradient-based TENT on full-precision 32-bit ViT, while achieving an up to 24-fold memory reduction on ImageNet-C. The source code will be released.
@@ 1 Introduction
Table 1: Comparison w.r.t. our Forward-Only Adaptation vs. prior gradient-based Test-Time Adaptation (TTA). The memory usage and accuracy are measured via ViT-Base and batch size 64 on ImageNet-C (level 5). The memory of 8-bit ViT is an ideal estimation by 0.25$\times$ memory of 32-bit ViT per Liu et al. (2021b).
Prior TTA Forward-Only Adaptation (ours)
@@
Gradient-free methods learn from test data by adapting the statistics in batch normalization layers (Schneider et al., 2020; Khurana et al., 2021; Lim et al., 2023), correcting the output probabilities (Boudiaf et al., 2022), or adjusting the classifier (Iwasawa & Matsuo, 2021), etc. These methods, which avoid backpropagation and do not alter the original model weights, inherently reduce the risk of forgetting on source domain. However, their limited learning capacity, primarily stemming from the constraint of not updating the core model parameters, may lead to suboptimal performance on out-of-distribution test data (as shown in Table 2).
@@
Firstly, deep models are usually deployed on various edge devices, e.g., smartphones and embedded systems. Unlike high-performance GPUs, these devices typically possess limited computational power and memory capacity, insufficient for the intensive computations required by TTA, which often needs one or multiple rounds of backpropagation for each test sample (Wang et al., 2021; Zhang et al., 2022).
Secondly, for resource or efficiency considerations, deep models often undergo quantization before deployment – a process of reducing precision, e.g., from 32-bit to 8-bit. However, the non-differentiability of the discrete quantizer would result in vanishing gradients when propagated through multiple layers (Louizos et al., 2019). This makes deployed models incapable of supporting backpropagation operations, which are essential for prior TTA methods.
@@
To address the above issues, we propose a test-time Forward-Only Adaptation (FOA) method. To be specific, we seek to explore a backpropagation-free optimizer, called covariance matrix adaptation (CMA) evolution strategy (Hansen, 2016), for online test-time model adaptation. However, naively applying CMA in the TTA setting is infeasible, as it is hard for CMA to handle ultra-high dimensional optimization problems (e.g., deep model training) and it relies on supervised learning signals. Therefore, we propose to solely update a newly inserted prompt (as the model’s input, shown in Figure 1) at test time to reduce the dimension of solution space and meanwhile avoid altering model weights. Then, to make CMA work stably without supervised signals, we devise a novel unsupervised fitness function to evaluate candidate solutions, which comprises both model prediction entropy and the activation statistics discrepancy between out-of-distribution (OOD) testing samples and source in-distribution (ID) samples. Here, only a small number of ID samples is needed for source statistics estimation, i.e., 32 samples are sufficient for ImageNet (see Figure 2 (c)). Moreover, to further boost the adaptation performance, we devise a forward-only back-to-source activation shifting mechanism to directly adjust the activations of OOD testing samples, along with a dynamically updated shifting direction from the OOD testing domain to the ID source domain.
Main Novelty and Contributions. 1) We introduce a novel and practical paradigm to TTA, termed forward-only adaptation. This paradigm operates without depending on backpropagation and avoids modification to the model weights, significantly broadening the real-world applicability of TTA across various contexts, including smartphones, FPGAs, and quantized models. 2) We achieve the goal of forward-only adaptation by prompt adaptation and activation shifting, where we design a new fitness function that guarantees stable prompt learning using a covariance matrix adaptation-based optimizer under the online unsupervised setting, and efficiently align the sample’s activations in the testing domain with the source training domain. 3) Extensive experiments on four benchmarks and full precision/quantized models verify our effectiveness. Our method on 8-bit quantized ViT outperforms gradient-based TENT on full-precision 32-bit ViT, with up to 24-fold run-time memory reduction.
@@ 2 Preliminary and Problem Statement
Vision Transformer (ViT) (Dosovitskiy et al., 2021). In this paper, we mainly focus on transformer-based vision models that are widely used in practice and are also hardware-friendly. We first revisit ViT here for the presentation convenience of our method. Formally, for a plain ViT $f_{\Theta}(\cdot)$ with $N$ layers, let ${\bf E}_{i}=\{{\bf e}_{i}^{j},j\in\mathbb{N},0\leq j\leq m\}$ be the patch embeddings as the input of the $(i+1)$-th layer $L_{i+1}$, where $m$ is the number of image patches and ${\bf e}_{i}^{0}$ denote an extra learnable classification token ([CLS]) of the $i$-th layer $L_{i}$, the whole ViT is formulated as:
@@ 3 Approach
In this paper, we propose a novel test-time Forward-Only Adaptation (FOA) method, which is also model updating-free, to boost the practicality of test-time adaptation in various real-world scenarios. From Figure 1, FOA conducts adaptation on both the input level and the output feature level. 1) Input level: FOA inserts a new prompt as the model’s input, and then solely updates this prompt online for out-of-distribution (OOD) generalization, employing a derivative-free optimizer coupled with a specially designed unsupervised fitness function (c.f. Section 3.1). 2) Output feature level: a back-to-source activation shifting strategy seeks to further boost adaptation, which directly refines the activation features of the final layer, by aligning them from the OOD domain back to the source in-distribution (ID) domain (c.f. Section 3.2). We summarize the pseudo-code of FOA in Algorithm 1.
@@ 3.1 Forward-Only Prompt Adaptation
Unlike prior TTA methods that update model weights using backpropagation, we aim to achieve the goal of test-time out-of-distribution generalization in a backpropagation-free manner. To this end, we explore a derivative-free optimizer for TTA, namely covariance matrix adaptation (CMA) evolution strategy (Hansen, 2016). However, naively applying CMA to our TTA context is infeasible, the reasons are due to: 1) For TTA, the model parameters needing update are high-dimensional (even for methods like TENT (Wang et al., 2021) that only updates the affine parameters of normalization layers), since the deep models are often with millions of parameters. This makes CMA intractable for direct deep model adaptation. 2) Conventional CMA methods rely on supervised offline learning, i.e., using ground-truth labels to assess the candidate solutions. In contrast, TTA operates without ground-truth labels and typically in an online setting, rendering conventional CMA methods inapplicable. We empirically illustrated these issues in Table 9.
@@
Algorithm 1 Forward-Only Adaptation (FOA).
@@
CMA-Based Prompt Adaptation. Inspired by that continuous prompt learning has been shown pretty effective in the field of natural language processing, we add prompt embeddings at the beginning of the model’s input (i.e., the first layer) for test-time updating, while keeping all other model parameters frozen. In this way, the dimension of learnable model parameters shall be significantly reduced and thus is compatible with CMA optimization. Formally, given a test sample ${\bf x}\sim Q({\bf x})$ and a ViT model $f_{\Theta}(\cdot)=\texttt{Head}(L_{i}(\cdot))$, our goal is to find an optimal prompt ${\bf p}^{*}$:
@@ 3.2 Back-to-Source Activation Shifting
where ${\bf d}$ is a shifting direction and $\gamma$ is a step size. We define ${\bf d}$ as the vector extending from the center of out-of-distribution testing features to the center of source in-distribution features. In our online TTA setting, with the increase of testing samples, the center of testing features shall dynamically change. Thus, we online update the shifting direction ${\bf d}$ by:
@@
where ${\bm{\mu}}_{N}^{S}$ is the mean of the $N$-th final layer CLS feature ${\bf e}_{N}^{0}$ and calculated over source in-distribution samples ${\mathcal{D}}_{S}$ (the same one used in Eqn. (3.1)). ${\bm{\mu}}_{N}(t)$ is the approximation of the overall test set statistics by exponential moving averages of statistics computed on the sequentially arrived test samples. We define the mean estimate of the ${\bf e}_{N}^{0}$ in iteration $t$ (the $t$-th batch ${\mathcal{X}}_{t}$) as:
@@ 4 Experiments
Implementation Details. We set the number of prompt embeddings $N_{p}$ to 3 and initialize prompts with uniform initialization. We set the batch size ($BS$) to 64 by following TENT and SAR for fair comparisons. The population size $K$ is set to $28=4+3\times\log(prompt~{}~{}dim)$ by following (Hansen, 2016) and $\lambda$ in Eqn. (3.1) is set to 0.4×$\times BS/64$ on ImageNet-C/V2/Sketch, and 0.2×$\times BS/64$ on ImageNet-R to balance the magnitude of two losses. We use the validation set of ImageNet-1K to estimate source ID statistics. The step size $\gamma$ in Eqn. (7) is set to 1. The effect of each hyperparameter is investigated in Section 4.3 and Appendix C. More details of compared methods are put in Appendix B.2.
@@
Table 2: Comparisons with SOTA methods on ImageNet-C (severity level 5) with ViT regarding Accuracy (%). BP is short for backward propagation and the bold number indicates the best result. We only report average ECE (%$\downarrow$) here and put detailed ECEs in Appendix.
@@ 4.1 Results on Full Precision Models
In this section, we compare our FOA with existing state-of-the-art TTA methods on the full precision ViT-Base model. From the results on ImageNet-C in Table 2, we have the following observations. 1) Our FOA achieves the best average accuracy and ECE over 15 different corruption types, suggesting our effectiveness. 2) Compared with NoAdapt, gradient-free methods LAME and T3A obtain slight performance gain or perform even worse, as they do not update core model weights and thus may suffer from limited learning capacity. Here, LAME performs worse than NoAdapt, because it only adjusts the model output logits and is not very effective when the OOD test sample stream does not suffer from prior label shifts, which is consistent with the results reported by LAME itself. 3) Compared with LAME and T3A, gradient-based methods (TENT, CoTTA and SAR) explicitly modify model parameters by optimizing unsupervised/self-supervised losses, and thus achieve much better performance, e.g., the average accuracy of 56.9% (T3A) vs. 62.7% (SAR). 4) Without using any back-propagation, our FOA outperforms gradient-based SAR with 3.6% average accuracy and 3.8% average ECE gains, demonstrating our superiority in deploying to lightweight devices (e.g., smartphones and FPGAs) and quantized models. 5) FOA achieves much lower average ECE compared with BP-based methods, e.g., 18.5% (TENT) vs. 3.2% (FOA). This mainly benefits from our activation discrepancy regularization (in Eqn. (3.1)), which alleviates the error accumulation issue of prior methods that may employ imprecise pseudo labels or entropy for learning. At last, from the results on ImageNet-R/V2/Sketch in Table 3, our FOA achieves the best or comparable performance w.r.t. both accuracy and ECE, further suggesting our effectiveness.
@@ 4.2 Results on Quantized Models
In practical applications, deep models are often deployed on edge devices with efficiency considerations, undergoing a process known as quantization. These devices, constrained by limited resources, typically do not support backward propagation due to its high memory and computational demands. Consequently, traditional gradient-based TTA methods like TENT, CoTTA, and SAR are not viable in such settings. In contrast, our FOA is adaptable to these quantized models. We demonstrate this by applying FOA to quantized ViT models and benchmarking it against T3A. As indicated in Table 4, FOA outperforms T3A significantly in terms of both accuracy and ECE on 8-bit and 6-bit models. Notably, FOA with an 8-bit ViT surpasses the performance of the gradient-based TENT method using a full precision 32-bit ViT on ImageNet-C, achieving 63.5% accuracy (our FOA, 8-bit) vs. 59.6% (TENT, 32-bit). These results collectively underscore the superiority of our FOA in such quantized model deployment scenarios.
@@ 4.3 Ablation Studies
Effectiveness of Components in FOA. In our FOA, we mentioned that naively applying CMA with entropy minimization in TTA is infeasible, and thus we propose an activation distribution discrepancy-based fitness function to guide the stable learning of CMA and an activation shifting scheme to boost the adaptation performance. We ablate them in Table 5. Firstly, CMA with Entropy fitness performs poorer than “NoAdapt”, which verifies the necessity for devising a new fitness function. Secondly, our Activation Discrepancy fitness works well to provide stable learning signals for CMA and improves the adaptation accuracy on ImageNet-C from 55.5% to 63.4%. Thirdly, even only with the Activation Shifting scheme, it also improves the accuracy from 55.5% to 59.1%, suggesting its effectiveness. Lastly, by incorporating Entropy and Activation Discrepancy as a whole fitness function and with the Activation Shifting scheme, our FOA achieves the best performance, i.e., 66.3% average accuracy and 3.2% average ECE on ImageNet-C.
@@
Effects of #Samples for Calculating $\{\bm{\mu}_{i}^{S},{\bm{\sigma}}_{i}^{S}\}_{i=0}^{N}$. As described in Section 3.1, the calculation of source training statistics $\{\bm{\mu}_{i}^{S},{\bm{\sigma}}_{i}^{S}\}_{i=0}^{N}$ involves a small set of unlabeled in-distribution samples, which can be collected via existing OOD detection techniques (Berger et al., 2021) or directly using the training samples. Here, we investigate the effect of #samples needed, selected from {16, 32, 64, 100, 200, 400, 800, 1600}. From Figure 2 (c), our FOA consistently achieves stable performance when #samples greater than 32, regarding both the accuracy and ECE. These results show that our FOA does not need to collect too many in-distribution samples, which are easy to obtain in practice.

### v2

@@ Abstract
Test-time adaptation has proven effective in adapting a given trained model to unseen test samples with potential distribution shifts. However, in real-world scenarios, models are usually deployed on resource-limited devices, e.g., FPGAs, and are often quantized and hard-coded with non-modifiable parameters for acceleration. In light of this, existing methods are often infeasible since they heavily depend on computation-intensive backpropagation for model updating that may be not supported. To address this, we propose a test-time Forward-Optimization Adaptation (FOA) method. In FOA, we seek to solely learn a newly added prompt (as model’s input) via a derivative-free covariance matrix adaptation evolution strategy. To make this strategy work stably under our online unsupervised setting, we devise a novel fitness function by measuring test-training statistic discrepancy and model prediction entropy. Moreover, we design an activation shifting scheme that directly tunes the model activations for shifted test samples, making them align with the source training domain, thereby further enhancing adaptation performance. Without using any backpropagation and altering model weights, FOA runs on quantized 8-bit ViT outperforms gradient-based TENT on full-precision 32-bit ViT, while achieving an up to 24-fold memory reduction on ImageNet-C. Code: https://github.com/mr-eggplant/FOA.
@@ 1 Introduction
Table 1: Comparison w.r.t. prior gradient-based Test-Time Adaptation (TTA) vs. our Forward-Optimization Adaptation. The memory usage and accuracy are measured via ViT-Base and batch size 64 on ImageNet-C (level 5). The memory of 8-bit ViT is an ideal estimation by 0.25$\times$ memory of 32-bit ViT per Liu et al. (2021b).
Prior TTA Forward-Optimization Adaptation
@@
Gradient-free methods learn from test data by adapting the statistics in batch normalization layers (Schneider et al., 2020; Khurana et al., 2021; Lim et al., 2023), correcting the output probabilities (Boudiaf et al., 2022), or adjusting the classifier (Iwasawa & Matsuo, 2021), etc. These methods, which avoid backpropagation and do not alter the original model weights, inherently reduce the risk of forgetting on source domain. However, their limited learning capacity, primarily stemming from the constraint of not explicitly exploiting the model feedback regarding given test samples to facilitate optimization with learnable parameters, may lead to suboptimal performance on out-of-distribution test data (as shown in Table 2).
@@
Firstly, deep models are usually deployed on various edge devices, e.g., smartphones, and embedded systems. Unlike high-performance GPUs, these devices typically possess limited computational power and memory capacity, insufficient for the intensive computations required by TTA, which often requires one or multiple rounds of backpropagation for each test sample (Wang et al., 2021; Zhang et al., 2022).
Secondly, for resource or efficiency considerations, deep models often undergo quantization before deployment – a process of reducing precision, e.g., from 32-bit to 8-bit. However, the non-differentiability of the discrete quantizer would result in vanishing gradients when propagated through multiple layers (Louizos et al., 2019). This makes the deployed models incapable of supporting backpropagation operations, which are essential for prior TTA methods.
@@
To address the above issues, we propose a test-time Forward-Optimization Adaptation (FOA) method. Specifically, we seek to explore a backpropagation-free optimizer, called covariance matrix adaptation (CMA) evolution strategy (Hansen, 2016), for online test-time model adaptation. However, naively applying CMA in the TTA setting is infeasible, as it is hard for CMA to handle ultra-high dimensional optimization problems (e.g., deep model training) and it relies on supervised learning signals. Therefore, we propose to solely update a newly inserted prompt (as the model’s input, shown in Figure 1) at test time to reduce the dimension of solution space and meanwhile avoid altering model weights. Then, to make CMA work stably without supervised signals, we devise a novel unsupervised fitness function to evaluate candidate solutions, which comprises both model prediction entropy and the activation statistics discrepancy between out-of-distribution (OOD) testing samples and source in-distribution (ID) samples. Here, only a small number of ID samples is needed for source statistics estimation, i.e., 32 samples are sufficient for ImageNet (see Figure 2 (c)). Moreover, to further boost adaptation performance, we devise a forward-only back-to-source activation shifting mechanism to directly adjust the activations of OOD testing samples, along with a dynamically updated shifting direction from the OOD testing domain to the ID source domain.
Main Contributions. 1) We introduce a novel and practical paradigm to TTA, termed forward-optimization adaptation. This paradigm operates without depending on backpropagation and avoids modification to the model weights, significantly broadening the real-world applicability of TTA across various contexts, including smartphones, FPGAs, and quantized models. 2) We achieve the goal of forward-only adaptation by prompt adaptation and activation shifting, where we design a new fitness function that guarantees stable prompt learning using a covariance matrix adaptation-based optimizer under the online unsupervised setting, and efficiently align the sample’s activations in the testing domain with the source training domain. 3) Extensive experiments on four benchmarks and full precision/quantized models verify our effectiveness. Our method on 8-bit quantized ViT outperforms gradient-based TENT on full-precision 32-bit ViT, with up to 24-fold run-time memory reduction.
@@ 2 Preliminary and Problem Statement
Vision Transformer (ViT) (Dosovitskiy et al., 2021). In this paper, we focus mainly on transformer-based vision models that are widely used in practice and are also hardware-friendly. We first revisit ViT here for the presentation convenience of our method. Formally, for a plain ViT $f_{\Theta}(\cdot)$ with $N$ layers, let ${\bf E}_{i}=\{{\bf e}_{i}^{j},j\in\mathbb{N},0\leq j\leq m\}$ be the patch embeddings as the input of the $(i+1)$-th layer $L_{i+1}$, where $m$ is the number of image patches and ${\bf e}_{i}^{0}$ denote an extra learnable classification token ([CLS]) of the $i$-th layer $L_{i}$, the whole ViT is formulated as:
@@ 3 Approach
In this paper, we propose a novel test-time Forward-Optimization Adaptation (FOA) method, which is also model updating-free, to boost the practicality of test-time adaptation in various real-world scenarios. From Figure 1, FOA conducts adaptation on both the input level and the output feature level. 1) Input level: FOA inserts a new prompt as the model’s input, and then solely updates this prompt online for out-of-distribution (OOD) generalization, employing a derivative-free optimizer coupled with a specially designed unsupervised fitness function (c.f. Section 3.1). 2) Output feature level: a back-to-source activation shifting strategy seeks to further boost adaptation, which directly refines the activation features of the final layer, by aligning them from the OOD domain back to the source in-distribution (ID) domain (c.f. Section 3.2). We summarize the pseudo-code of FOA in Algorithm 1.
@@ 3.1 Forward-Only Prompt Adaptation
Unlike prior TTA methods that update model weights using backpropagation, we aim to achieve the goal of test-time out-of-distribution generalization in a backpropagation-free manner. To this end, we explore a derivative-free optimizer for TTA, namely covariance matrix adaptation (CMA) evolution strategy (Hansen, 2016). However, naively applying CMA to our TTA context is infeasible, the reasons are due to: 1) For TTA, the model parameters needing update are high-dimensional (even for methods like TENT (Wang et al., 2021) that only updates the affine parameters of normalization layers), since the deep models are often with millions of parameters. This makes CMA intractable for direct deep model adaptation. 2) Conventional CMA methods rely on supervised offline learning, i.e., using ground-truth labels to assess candidate solutions. In contrast, TTA operates without ground-truth labels and typically in an online setting, rendering conventional CMA methods inapplicable. We empirically illustrated these issues in Table 9.
@@
Algorithm 1 Forward-Optimization Adaptation (FOA).
@@
CMA-Based Prompt Adaptation. Inspired by the demonstrated effectiveness of continuous prompt learning in the field of deep model fine-tuning (Jia et al., 2022; Bahng et al., 2022), we add new prompt embeddings at the beginning of the model input (i.e., before the first transformer layer) for test-time updating, while keeping all other model parameters frozen. In this way, the dimension of learnable model parameters shall be significantly reduced and thus is compatible with CMA optimization. Formally, given a test sample ${\bf x}\sim Q({\bf x})$ and a ViT model $f_{\Theta}(\cdot)=\texttt{Head}(L_{i}(\cdot))$, our goal is to find an optimal prompt ${\bf p}^{*}$:
@@ 3.2 Back-to-Source Activation Shifting
where ${\bf d}$ is a shifting direction and $\gamma$ is a step size. We define ${\bf d}$ as the vector extending from the center of out-of-distribution testing features to the center of source in-distribution features. In our online TTA setting, with the increase of testing samples, the center of testing features shall dynamically change. Thus, we update the shifting direction ${\bf d}$ online by:
@@
where ${\bm{\mu}}_{N}^{S}$ is the mean of the $N$-th final layer CLS feature ${\bf e}_{N}^{0}$ and calculated over source in-distribution samples ${\mathcal{D}}_{S}$ (the same one used in Eqn. (3.1)). ${\bm{\mu}}_{N}(t)$ is the approximation of the overall test set statistics by exponential moving averages of statistics computed on sequentially arrived test samples. We define the mean estimate of the ${\bf e}_{N}^{0}$ in iteration $t$ (the $t$-th batch ${\mathcal{X}}_{t}$) as:
@@ 4 Experiments
Implementation Details. We set the number of prompt embeddings $N_{p}$ to 3 and initialize prompts with uniform initialization. We set the batch size ($BS$) to 64 by following TENT and SAR for fair comparisons. The population size $K$ is set to $28=4+3\times\log(prompt~{}~{}dim)$ by following (Hansen, 2016) and $\lambda$ in Eqn. (3.1) is set to 0.4×$\times BS/64$ on ImageNet-C/V2/Sketch, and 0.2×$\times BS/64$ on ImageNet-R to balance the magnitude of two losses. We use the validation set of ImageNet-1K to estimate source ID statistics. The step size $\gamma$ in Eqn. (7) is set to 1.0, aiming to exactly align the overall center of testing and training features. The effect of each hyperparameter is investigated in Section 4.3 and Appendix C. More implementation details of compared methods are put in Appendix B.2.
@@
Table 2: Comparisons with SOTA methods on ImageNet-C (severity level 5) with ViT regarding Accuracy (%). BP is short for backward propagation and the bold number indicates the best result. We only report average ECE (%$\downarrow$) here and put detailed ECEs in Appendix D.
@@ 4.1 Results on Full Precision Models
In this section, we compare our FOA with existing state-of-the-art TTA methods on the full precision ViT-Base model. From the results on ImageNet-C in Table 2, we have the following observations. 1) Our FOA achieves the best average accuracy and ECE over 15 different corruption types, suggesting our effectiveness. 2) Compared with NoAdapt, gradient-free methods LAME and T3A obtain slight performance gains or perform even worse, as they do not update core model weights and thus may suffer from limited learning capacity. Here, LAME performs worse than NoAdapt, because it only adjusts the model output logits and is not very effective when the OOD test sample stream does not suffer from prior label shifts, which is consistent with the results reported by LAME itself. 3) Compared with LAME and T3A, gradient-based methods (TENT, CoTTA and SAR) explicitly modify model parameters by optimizing unsupervised/self-supervised losses, and thus achieve much better performance, e.g., the average accuracy of 56.9% (T3A) vs. 62.7% (SAR). 4) Without using any back-propagation, our FOA outperforms gradient-based SAR with 3.6% average accuracy and 3.8% average ECE gains, demonstrating our superiority in deploying to lightweight devices (e.g., smartphones and FPGAs) and quantized models. 5) FOA achieves much lower average ECE compared with BP-based methods, e.g., 18.5% (TENT) vs. 3.2% (FOA). This mainly benefits from our activation discrepancy regularization (in Eqn. (3.1)), which alleviates the error accumulation issue of prior methods that may employ imprecise pseudo labels or entropy for learning. At last, from the results on ImageNet-R/V2/Sketch in Table 3, our FOA achieves the best or comparable performance w.r.t. both accuracy and ECE, further suggesting our effectiveness.
@@ 4.2 Results on Quantized Models
In practical applications, deep models are often deployed on edge devices with efficiency considerations, undergoing a process known as quantization. These devices, constrained by limited resources, typically do not support backward propagation due to its high memory and computational demands. Consequently, traditional gradient-based TTA methods like TENT, CoTTA, and SAR are not viable in such settings. On the contrary, our FOA is adaptable to these quantized models. We demonstrate this by applying FOA to quantized ViT models and benchmarking it against T3A. As indicated in Table 4, FOA outperforms T3A significantly in terms of both accuracy and ECE on 8-bit and 6-bit models. Notably, FOA with an 8-bit ViT surpasses the performance of the gradient-based TENT method using a full precision 32-bit ViT on ImageNet-C, achieving 63.5% accuracy (our FOA, 8-bit) vs. 59.6% (TENT, 32-bit). These results collectively underscore the superiority of our FOA in such quantized model deployment scenarios.
@@ 4.3 Ablation Studies
Effectiveness of Components in FOA. In our FOA, we mentioned that naively applying CMA with entropy minimization in TTA is infeasible, and thus we propose an activation distribution discrepancy-based fitness function to guide the stable learning of CMA and an activation shifting scheme to boost the adaptation performance. We ablate them in Table 5. Firstly, CMA with Entropy fitness performs poorer than “NoAdapt”, which verifies the necessity of devising a new fitness function. Secondly, our Activation Discrepancy fitness works well to provide stable learning signals for CMA and improves the adaptation accuracy on ImageNet-C from 55.5% to 63.4%. Thirdly, even only with the Activation Shifting scheme, it also improves the accuracy from 55.5% to 59.1%, suggesting its effectiveness. Lastly, by incorporating Entropy and Activation Discrepancy as a whole fitness function and with the Activation Shifting scheme, our FOA achieves the best performance, i.e., 66.3% average accuracy and 3.2% average ECE on ImageNet-C.
@@
Effects of #Samples ($Q$) for Calculating $\{\bm{\mu}_{i}^{S},{\bm{\sigma}}_{i}^{S}\}_{i=0}^{N}$. As described in Section 3.1, the calculation of source training statistics $\{\bm{\mu}_{i}^{S},{\bm{\sigma}}_{i}^{S}\}_{i=0}^{N}$ involves a small set of unlabeled in-distribution samples, which can be collected via existing OOD detection techniques (Berger et al., 2021) or directly using the training samples. Here, we investigate the effect of #samples needed, selected from {16, 32, 64, 100, 200, 400, 800, 1600}. From Figure 2 (c), our FOA consistently achieves stable performance when #samples greater than 32, regarding both the accuracy and ECE. These results show that our FOA does not need to collect too many in-distribution samples, which are easy to obtain in practice.